A journal published a study of​ travelers' familiarity with ecolabels used by hotels. Adult travelers were shown a list of 6 different​ ecolabels, and​ asked, "How familiar are you with this​ ecolabel?" Suppose the response is measured on a continuous scale from 10​ (not familiar at​ all) to 50​ (very familiar), and the mean and standard deviation for ecolabel E are 44 and 1.5​, respectively. Assume the distribution of responses is approximately normal.
a. Find the probability that a response to ecolabel E exceeds 43.
The probability is

Answer by stanbon(75887)

1:17): A journal published a study of​ travelers' familiarity with ecolabels used by hotels. Adult travelers were shown a list of 6 different​ ecolabels, and​ asked, "How familiar are you with this​ ecolabel?" Suppose the response is measured on a continuous scale from 10​ (not familiar at​ all) to 50​ (very familiar), and the mean and standard deviation for ecolabel E are 44 and 1.5​, respectively. Assume the distribution of responses is approximately normal.
a. Find the probability that a response to ecolabel E exceeds 43.
z(43) = (43-44)/1.5 = -1/1.5 = -2/3
----
Ans:: P(x > 43) = P(z > -2/3) = 1-normalcdf(-100,-2/3) = 0.7475
Cheers,
Stan H.
----------
The probability is 0.7475